OpenStudy (laddiusmaximus):

im confused on this step 3 f(2) + f(2) * 6 =36 becomes 9f(2)=36 how did this happen?

OpenStudy (lgbasallote):

do you agree that f(2) * 6 = 6f(2) ?

OpenStudy (laddiusmaximus):

yes

OpenStudy (lgbasallote):

so you have 3f(2) + 6f(2) = 36 now think of f(2) as x do you know what 3x + 6x equals?

OpenStudy (laddiusmaximus):

9x

OpenStudy (lgbasallote):

right. and that is why 3f(2) + 6f(2) = 9f(2) make sense?

OpenStudy (laddiusmaximus):

now it does yes. thank you

OpenStudy (lgbasallote):

welcome